OpenStudy (anonymous):

(6x^4y^7)^3 simplify!

OpenStudy (anonymous):

please show steps to simplifying :)

OpenStudy (anonymous):

damn

OpenStudy (anonymous):

lookin joocy

OpenStudy (anonymous):

thick. solid. tight mmmmm

OpenStudy (anonymous):

\[(6x^4y^7)^3\]\[(a^b)^c=a^{b\times c}\]so that means\[(6)^3(x^4)^3(y^7)^3~~\implies~~6^3(x^{4\times 3})(y^{7\times 3})\]you wanna do the rest? :)

OpenStudy (anonymous):

@xxfreshboy

OpenStudy (anonymous):

uhm would mean alot if you could help finsih it I suck at this stuff!

OpenStudy (anonymous):

Anyway sexy, when an exponent is being raised by another exponent, all you do is multiply. (You will multiply 4 and 3, and 7 and 3 in this case) You also have to remember the 6. Raise that to the third power. 6*6*6 = 216 216x^12y^21

OpenStudy (anonymous):